During practice, a basketball was tossed upward with a speed of 15.0 m/s. It is caught at the same distance above the ground. (a) How high did the ball rise? (b) How long did the ball remain in the air?
1
Expert's answer
2020-10-15T10:45:54-0400

Let us choose the origin, which coincides with the initial position of the ball. Then the equations of vertical motion are "y(t) = v_0 t - \frac{g t^2}{2}", "v(t) = v_0 - g t".

At highest point, the speed is zero, hence "v(t') = v_0 - g t'", from where the time to reach that point is "t' = \frac{v_0}{g}". Hence, the ball rose at height "y(t') = \frac{v_0^2}{g} - \frac{g}{2} \frac{v_0^2}{g^2} = \frac{v_0^2}{2 g} \approx 11.47 m".


To find the time the ball remains in the air, equate its y coordinate to zero, and solve for time:

"y(t) = v_0 t - \frac{g t^2}{2} = t(v_0 - \frac{g t}{2}) = 0", from where "t = 0" (initially the ball is at the origin) or "t = \frac{2 v_0}{g} \approx 3.1 s". Hence, the ball remains "3.1 s" in the air.